MKM International is seeking to purchase a new CNC machine in order to reduce costs. Two alternative machines are in consideration. Machine 1 costs $ 450 comma 000​, but yields a 15 percent savings over the current machine used. Machine 2 costs $ 950 comma 000​, but yields a 25 percent savings over the current machine used. In order to meet​ demand, the following forecasted cost information for the current machine is also provided.                                                                                   LOADING... Year Project Cost 1 950 comma 000 2 1 comma 350 comma 000 3 1 comma 450 comma 000 4 1 comma 550 comma 000 5 2 comma 400 comma 000 a. Based on the NPV of the cash flows for these 5​ years, which machine should MKM International​ purchase? Assume a discount rate of 13 percent. Assuming a discount rate of 13 ​percent, MKM International should purchase ▼ because the NPV of machine 1 is ​$ nothing and the NPV of machine 2 is ​$ nothing. ​(Enter your responses rounded to the nearest whole​ number.) b. If MKM International lowered its required discount rate to 7 ​percent, what machine would it​ purchase? Assuming a discount rate of 7 ​percent, MKM International should purchase ▼ machine 1 machine 2 because the NPV of machine 1 is ​$ nothing and the NPV of machine 2 is ​$ nothing. ​(Enter your responses rounded to the nearest whole​ number.)
Answer #1

Answer in MS Excel Sheet with step by step Solution:

Machine 1
Statement Type Year Amount in USD Factor for discount rate @ 13% Present Value in USD
Original Cost Cash Outflow 0 -450000 1 -450000
Project Cost Cash Outflow 1 -950000 0.884955752 -840707.9646
Project Cost Cash Outflow 2 -1350000 0.783146683 -1057248.023
Project Cost Cash Outflow 3 -1450000 0.693050162 -1004922.735
Project Cost Cash Outflow 4 -1550000 0.613318728 -950644.0279
Project Cost Cash Outflow 5 -2400000 0.542759936 -1302623.846
Saving Cash Inflow 5 67500 0.542759936 36636.29568
Net Present Value or NPV = -5569510.301 (Outflow)= -5569510
Machine 2
Statement Type Year Amount in USD Factor for discount rate @ 13% Present Value in USD
Original Cost Cash Outflow 0 -950000 1 -950000
Project Cost Cash Outflow 1 -950000 0.884955752 -840707.9646
Project Cost Cash Outflow 2 -1350000 0.783146683 -1057248.023
Project Cost Cash Outflow 3 -1450000 0.693050162 -1004922.735
Project Cost Cash Outflow 4 -1550000 0.613318728 -950644.0279
Project Cost Cash Outflow 5 -2400000 0.542759936 -1302623.846
Saving Cash Inflow 5 237500 0.542759936 128905.4848
Net Present Value or NPV = -5977241.112 (Outflow) = -5977241

Interpretation: as per above Table and calculation, using discount rate of 13%, we find that NPV for using Machine A is Higher than that of Machine B, Thus, we will select Machine A for this project purpose, NPV Value for Machine A and Machine B written in the MS Excel Sheet,

Formula Presented for Reference:

1 B C D E F G
2 Machine 1
3 Statement Type Year Amount in USD Factor for discount rate @ 13% Present Value in USD
4 Original Cost Cash Outflow 0 -450000 =1/(1+0.13)^D4 =E4*F4
5 Project Cost Cash Outflow 1 -950000 =1/(1+0.13)^D5 =E5*F5
6 Project Cost Cash Outflow 2 -1350000 =1/(1+0.13)^D6 =E6*F6
7 Project Cost Cash Outflow 3 -1450000 =1/(1+0.13)^D7 =E7*F7
8 Project Cost Cash Outflow 4 -1550000 =1/(1+0.13)^D8 =E8*F8
9 Project Cost Cash Outflow 5 -2400000 =1/(1+0.13)^D9 =E9*F9
10 Saving Cash Inflow 5 =-0.15*E4 =1/(1+0.13)^D10 =E10*F10
11 Net Present Value or NPV = =SUM(G4:G10)
12
13 Machine 2
14 Statement Type Year Amount in USD Factor for discount rate @ 13% Present Value in USD
15 Original Cost Cash Outflow 0 -950000 =1/(1+0.13)^D15 =E15*F15
16 Project Cost Cash Outflow 1 -950000 =1/(1+0.13)^D16 =E16*F16
17 Project Cost Cash Outflow 2 -1350000 =1/(1+0.13)^D17 =E17*F17
18 Project Cost Cash Outflow 3 -1450000 =1/(1+0.13)^D18 =E18*F18
19 Project Cost Cash Outflow 4 -1550000 =1/(1+0.13)^D19 =E19*F19
20 Project Cost Cash Outflow 5 -2400000 =1/(1+0.13)^D20 =E20*F20
21 Saving Cash Inflow 5 =-0.25*E15 =1/(1+0.13)^D21 =E21*F21
22 Net Present Value or NPV = =SUM(G15:G21)
Machine 1
Statement Type Year Amount in USD Factor for discount rate @ 7% Present Value in USD
Original Cost Cash Outflow 0 -450000 1 -450000
Project Cost Cash Outflow 1 -950000 0.934579439 -887850.4673
Project Cost Cash Outflow 2 -1350000 0.873438728 -1179142.283
Project Cost Cash Outflow 3 -1450000 0.816297877 -1183631.921
Project Cost Cash Outflow 4 -1550000 0.762895212 -1182487.579
Project Cost Cash Outflow 5 -2400000 0.712986179 -1711166.831
Saving Cash Inflow 5 67500 0.712986179 48126.56712
Net Present Value or NPV = -6546152.514(Outflow) = -6546153
Machine 2
Statement Type Year Amount in USD Factor for discount rate @ 7% Present Value in USD
Original Cost Cash Outflow 0 -950000 1 -950000
Project Cost Cash Outflow 1 -950000 0.934579439 -887850.4673
Project Cost Cash Outflow 2 -1350000 0.873438728 -1179142.283
Project Cost Cash Outflow 3 -1450000 0.816297877 -1183631.921
Project Cost Cash Outflow 4 -1550000 0.762895212 -1182487.579
Project Cost Cash Outflow 5 -2400000 0.712986179 -1711166.831
Saving Cash Inflow 5 237500 0.712986179 169334.2176
Net Present Value or NPV = -6924944.864(Outflow) = -6924945

Interpretation: as per above Table and calculation, using discount rate of 7%, we find that NPV for using Machine A is Higher than that of Machine B, Thus, we will select Machine A for this project purpose, NPV Value for Machine A and Machine B written in the MS Excel Sheet,

Formula Presented for Reference:

1 B C D E F G
2 Machine 1
3 Statement Type Year Amount in USD Factor for discount rate @ 7% Present Value in USD
4 Original Cost Cash Outflow 0 -450000 =1/(1+0.07)^D4 =E4*F4
5 Project Cost Cash Outflow 1 -950000 =1/(1+0.07)^D5 =E5*F5
6 Project Cost Cash Outflow 2 -1350000 =1/(1+0.07)^D6 =E6*F6
7 Project Cost Cash Outflow 3 -1450000 =1/(1+0.07)^D7 =E7*F7
8 Project Cost Cash Outflow 4 -1550000 =1/(1+0.07)^D8 =E8*F8
9 Project Cost Cash Outflow 5 -2400000 =1/(1+0.07)^D9 =E9*F9
10 Saving Cash Inflow 5 =-0.15*E4 =1/(1+0.07)^D10 =E10*F10
11 Net Present Value or NPV = =SUM(G4:G10)
12
13 Machine 2
14 Statement Type Year Amount in USD Factor for discount rate @ 7% Present Value in USD
15 Original Cost Cash Outflow 0 -950000 =1/(1+0.07)^D15 =E15*F15
16 Project Cost Cash Outflow 1 -950000 =1/(1+0.07)^D16 =E16*F16
17 Project Cost Cash Outflow 2 -1350000 =1/(1+0.07)^D17 =E17*F17
18 Project Cost Cash Outflow 3 -1450000 =1/(1+0.07)^D18 =E18*F18
19 Project Cost Cash Outflow 4 -1550000 =1/(1+0.07)^D19 =E19*F19
20 Project Cost Cash Outflow 5 -2400000 =1/(1+0.07)^D20 =E20*F20
21 Saving Cash Inflow 5 =-0.25*E15 =1/(1+0.07)^D21 =E21*F21
22 Net Present Value or NPV = =SUM(G15:G21)